OpenStudy (anonymous):

Solve for x e^(2x + 5) - 5 = 0 Please show steps

myininaya (myininaya):

add 5 to both sides

myininaya (myininaya):

write 5 as \[e^{\ln(5)}\]

myininaya (myininaya):

so we have \[e^{2x+5}=e^{\ln(5)}\]

OpenStudy (anonymous):

e^(2x+5) = 5 (2x + 5) ln e = ln5 2x + 5 = ln5/lne x = ((ln 5/ln e) - 5) / 2

myininaya (myininaya):

since the bases are the same then the exponents are the same

myininaya (myininaya):

\[2x+5=\ln(5) => 2x=\ln(5)-5 => x=\frac{\ln(5)-5}{2}\]